A cylinder of metal with a mass of 23.6546 g is submerged in a graduated cylinder containing 25.20 mL of water. The water level rises to 29.90 mL. Which of the following densities is recorded to the correct number of significant digits?

5.0 g/mL

5.03 g/mL

5.032 g/mL

5.033 g/mL

5.0329 g/mL
Answer #1

density = mass / volume

mass = 23.6546 g/mol

volume = 29.90 - 25.20 =4.70 mL

density = 23.6546 / 4.70

density = 5.0329 g/mL

answer = 5.0329 g/mL